OpenStudy (anonymous):

What potential difference is measured across a 17.4 Ω load resistor when it is connected across a battery of emf 3.68 V and internal resistance 0.388 Ω?

OpenStudy (anonymous):

Are you sure that's the question

OpenStudy (anonymous):

yes

OpenStudy (shamim):

So given R=17.4 ohm E=3.68 volt r=0.388 ohm

OpenStudy (shamim):

We know I=E/(R+r)=?

OpenStudy (shamim):

Now E=V+Ir V=E-Ir=?

OpenStudy (shamim):

Response plz!!!!